OpenStudy (anonymous):

Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x^2 + 3x + 2 y = x^2 + 2x + 1

OpenStudy (anonymous):

do you mean find the roots or x-intercepts? i.e., solve 0=x^2+3x+2 0=x^2+2x+1

OpenStudy (anonymous):

Mackenzie kremer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh lol sorry

OpenStudy (anonymous):

asker? you there?